# biblio.bib

### v2

@@ 3.2 Algorithm description
3:Initialize $\lambda\leftarrow$ minimum improving ratio associated to $(\Pi,\Psi)$
@@
5:Initialize $\varepsilon\leftarrow\frac{\lambda}{100c_{1}\operatorname{log}n}$triangleright $c_{1}$ is a constant given by the proof of Theorem 3.7
@@
9:     Run MPX with $\rho=10\cdot c_{2}\cdot\varepsilon^{2}/(c\operatorname{log}^{2}n)$triangleright $d=O(\operatorname{log}n/\rho)$ by Lemma 3.6
@@
We now give a formal description of the algorithm given in Algorithm 1. It solves an LOP $(\Pi,\Psi)$ as defined in Definition 2.7. It takes as input the number of nodes $n$ and the description of the LOP $(\Pi,\Psi)$. In more detail, it gets the full description of $\Pi=(\mathcal{V}_{\operatorname{in}},\mathcal{V}_{\operatorname{out}},\mathcal{E}_{\operatorname{in}},\mathcal{E}_{\operatorname{out}},\mathcal{C})$, where $\mathcal{C}$ is an $(2r,\Delta)$-set of constraints over $(\mathcal{V}_{\operatorname{in}}\times\mathcal{V}_{\operatorname{out}},\mathcal{E}_{\operatorname{in}}\times\mathcal{E}_{\operatorname{out}})$, and the potential function $\Psi$ that assigns a non-negative real value to every element of $\mathcal{L}(\mathcal{V}_{\operatorname{in}}\times\mathcal{V}_{\operatorname{out}},\mathcal{E}_{\operatorname{in}}\times\mathcal{E}_{\operatorname{out}},r/2,\Delta)$.
@@
Consider any input graph $G=(V,E)$ and any labeling $\ell$ of $G$. For every centered graph $(H,v_{H})\in\mathcal{L}(\mathcal{V}_{\operatorname{in}}\times\mathcal{V}_{\operatorname{out}},\mathcal{E}_{\operatorname{in}}\times\mathcal{E}_{\operatorname{out}},r,\Delta)$ that does not belong to $\mathcal{C}$, such that $H$ is a subgraph of $G$, there is a way to reassign the labels to $v_{H}$ and its incident half-edges so that the potential $\operatorname{Pot}(G,\ell)$ decreases. Let $\ell^{\prime}$ be the new labeling: ($(\{v_{H}\},\ell^{\prime})$ is a minimal improving set with some positive improving ratio. Furthermore, the value $\operatorname{IR}(\{v_{H}\},\ell,\ell^{\prime})$ is completely determined by the radius-$2r$ neighborhood of $v_{H}$. Let $\lambda$ be the minimum possible improving ratio of the potential over any centered graph $(H,v_{H})\in\mathcal{L}(\mathcal{V}_{\operatorname{in}}\times\mathcal{V}_{\operatorname{out}},\mathcal{E}_{\operatorname{in}}\times\mathcal{E}_{\operatorname{out}},2r,\Delta)$ that does not belong to $\mathcal{C}$. We say that $\lambda$ is the minimum improving ratio associated to $(\Pi,\Psi)$ (which is initialized in Line 3). Similarly, let $\Lambda$ be the maximum possible decrease of the potential over any centered graph $(H,v_{H})\in\mathcal{L}(\mathcal{V}_{\operatorname{in}}\times\mathcal{V}_{\operatorname{out}},\mathcal{E}_{\operatorname{in}}\times\mathcal{E}_{\operatorname{out}},2r,\Delta)$ that does not belong to $\mathcal{C}$. We say that $\Lambda$ is the maximum improvement associated to $(\Pi,\Psi)$.
@@
The idea of the algorithm is as follows. Set $\varepsilon=\lambda/(100c_{1}\operatorname{log}(n))$, where $c_{1}$ is a small enough constant (Line 5), and set a minimum improving ratio to be $R=\lambda/4$ (Line 4). We start with an arbitrary labeling of the nodes and half-edges of $G$. Then, we have $\Theta(\operatorname{log}n)$ phases (from Line 8 to Line 26), where in each phase we run the MPX clustering algorithm (Line 9) with parameters $\rho=10\cdot c_{2}\cdot\varepsilon^{2}/(c\operatorname{log}^{2}n)=O(\operatorname{log}^{4}n)$ and $d=O(\operatorname{log}n/\rho)$, where $c,c_{2}$ are large enough constants that come from Lemmas 3.6 and 4.8, respectively: the bound on $d$ then holds by Lemma 3.6. Each run of MPX takes roughly $O(d)=O(\operatorname{log}^{5}n)$ rounds. We assume that each cluster $C$ given by the MPX algorithm comes with a leader node $v_{C}$, that can be the center of the cluster itself. This anyway can be obtained in $O(d)$ rounds after each MPX run. Each leader node looks at the subgraph $H$ of $G$ induced by the cluster $C$ it belongs to, and it computes a maximal sequence of $R$-improving sets $(A_{i},\ell_{i})_{i=1}^{k}$ in $H$ w.r.t. $\ell$. The sequence satisfies that $\mathcal{N}_{2r+1}[A_{i}]\subseteq C$ for all $i\in[k]$ (from Line 13 to Line 14), ensuring that the sequences of different clusters do not interfere with each other. Furthermore, we require the strong diameter of each set in the sequence to be at most $c_{3}\operatorname{log}n/\varepsilon$ (the choice of the large constant $c_{3}$ comes from Lemma 4.5). When we say that the sequence is maximal, we mean that in $H$ labeled by $\ell_{k}$ there is no minimal improving set $(A^{\prime},\ell^{\prime})$ such that $\operatorname{IR}(A^{\prime},\ell_{k},\ell^{\prime})>R$ and $\mathcal{N}_{2r+1}[A^{\prime}]\subseteq H$. Note that such a sequence must exist and terminate, as the potential of $H$ is always reduced after every relabeling (and cannot be reduced arbitrarily): furthermore, the sequence can be computed in $O(d)$ rounds by brute-force. For this the leader simply collects the topology and labeling of its entire cluster and brute forces every possible sequence in local computation. Then, $v_{C}$ adopts the labeling defined by $\ell_{k}$ and also broadcasts $\ell_{k}$ to every node of the cluster $C$ (Line 17). Every non-leader node just waits for the leader node to broadcast the new labeling (lines 20 and 21), and then it adopts the new labeling (Line 22). Then, the phase terminates and the nodes update the reference value of the improving ratio $R$ to $R+\lambda/(20c_{1}\operatorname{log}n)$ (Line 25). The algorithm terminates after $\Theta(\operatorname{log}n)$ phases, and it returns the labels of the nodes and of the incident half-edges (Line 27).
@@ 4 Analysis of the algorithm
For the rest of the section, we fix an LOP of interest $(\Pi,\Psi)$ and any input graph $G=(V,E)$ of $n$ nodes that is already $(\mathcal{V}_{\operatorname{in}}\times\mathcal{V}_{\operatorname{out}},\mathcal{E}_{\operatorname{in}}\times\mathcal{E}_{\operatorname{out}})$-labeled (hence it has an input and a – possibly invalid – output labeling). Let $\ell$ denote the labeling of $G$. Let $\lambda$ be the minimum possible improving ratio of the potential over any centered graph $(H,v_{H})\in\mathcal{L}(\mathcal{V}_{\operatorname{in}}\times\mathcal{V}_{\operatorname{out}},\mathcal{E}_{\operatorname{in}}\times\mathcal{E}_{\operatorname{out}},2r,\Delta)$ that does not belong to $\mathcal{C}$. We say that $\lambda$ is the minimum improving ratio associated to $(\Pi,\Psi)$. Similarly, let $\Lambda$ be the maximum possible decrease of the potential over any centered graph $(H,v_{H})\in\mathcal{L}(\mathcal{V}_{\operatorname{in}}\times\mathcal{V}_{\operatorname{out}},\mathcal{E}_{\operatorname{in}}\times\mathcal{E}_{\operatorname{out}},2r,\Delta)$ that does not belong to $\mathcal{C}$. We say that $\Lambda$ is the maximum improvement associated to $(\Pi,\Psi)$.
We remark that, being any LOP, all the values $r,\Delta,\lambda,\Lambda$ are constant values and will get lost in asymptotic notations. In the proofs of the lemmas in Section 4.1, we do keep track of a loose dependency in these parameters, which is polynomial. The reader might extract a $\text{poly}(\Delta^{r},\Lambda)\cdot\text{polylog}(n)$-time algorithm from this analysis; however, we remark that it is not in our interest to optimize dependencies from these parameters in this first work, which is why we do not keep explicit track of them.
We fix some standard notation, for the best possible labeling of a given set $A$, so we do not have to introduce these every time.
@@
For any subset of nodes $A\subseteq V$, we denote by $\ell^{\star}_{A}$ the best possible labeling, that is,
$\ell^{\star}_{A}=\arg\max_{x\in\mathscr{L}}\set{\operatorname{Imp}(A,\ell,x)}$
where $\mathscr{L}$ is the set of all possible labelings of $G$ such that $G$ is a $(\mathcal{V}_{\operatorname{in}}\times\mathcal{V}_{\operatorname{out}},\mathcal{E}_{\operatorname{in}}\times\mathcal{E}_{\operatorname{out}})$-labeled graph where the input labels are untouched.
The whole analysis of the algorithm relies on the proof of the following two lemmas (Lemmas 4.2 and 4.4), the last of which requires involved technical claims. The first lemma states that each phases terminates.
@@
For any phase $j$ of the algorithm, let $\ell_{j}$ be the labeling of the algorithm after finishing phase $j-1$ (Lines 1823) and $R_{j}$ be the value of $R$ at the end of phase $j-1$ (Line 25). Furthermore, let $C^{(j)}_{1},\ldots,C^{(j)}_{k}$ be the clusters that are the result of the MPX algorithm at phase $j$ (Line 9), Then, for each cluster $C^{(j)}_{i}$, any maximal sequence of $R_{j}$-improving sets is finite.
@@
So $B_{i}$ is the set of all nodes that have an edge that is “crossing different clusters” after running MPX at phase $i$. The next lemma shows that after phase $i$, all remaining improving sets of small diameter that have improving ratio $R_{i}$ (the value of $R$ initialized at the end of phase $i-1$ at Line 25) are close to all borders $B_{j}$ for all $j\leq i$.
Note that after the last phase the intersection of all border nodes will be empty with high probability. So this lemma implies that at the end of our algorithm, there do not exist any improving sets with small diameter and good improving ratio anymore.
@@
At the end of phase $i$ of the algorithm, all improving sets of diameter at most $O(\operatorname{log}n/\varepsilon)$ and improving ratio at least $R_{i}$ (the value of $R$ initialized at the end of phase $i-1$ at Line 25) are fully contained in the set
@@
The proof of Lemma 4.4 relies on the following two lemmas. The first one states that for any minimal improving set $A$ of some improving ratio $\beta$, for any node $v\in A$, we can find, within some radius-$O(\operatorname{log}n/\varepsilon)$ neighborhood of $v$, an improving set $A^{\prime}\subseteq A$ with improving ratio at least $\beta-\varepsilon$.
This lemma is what sparked the idea for this algorithm. Informally, using terminology from the introduction, let $X$ be the set of nodes flipped in a sequential flipping process that leads to a correct solution. Since the minimum improvement is $\lambda$ and every flip in the process makes an unhappy node happy, $X$ has improving ratio $\lambda$. The following lemma now says (in much more generic terms), that for every node $v\in X$, the $O(\operatorname{log}n/\varepsilon)$ neighborhood of $v$ already contains a $\lambda-\varepsilon$ improving set.
@@
Let $\ell$ be a labeling of $G=(V,E)$ and let $(S,\ell_{S})$ be a $\beta$-minimal improving set, for some $0<\beta\leq 1$. Then, for every $0<\varepsilon<\beta$ and any $v\in S$ there exists a large enough constant $c_{3}$ and a $k\leq c_{3}\operatorname{log}n/\varepsilon$ and a subset $A\subseteq S\cap\mathcal{N}_{k}[v]$, such that $\operatorname{IR}(A,\ell,\ell_{A})=\beta-\varepsilon$, where $\ell_{A}$ is equal to $\ell_{S}$ for all nodes in $A$ and their incident half-edges, while it is equal to $\ell$ for all other nodes and half-edges.
The proof of Lemma 4.5 is involved, and we defer to Section 4.1. The second lemma states that in a sequence of $\beta$-improving sets with small diameter, before any relabeling, around each node of the sequence there is an improving set $A$ with improving ratio slightly less than $\beta$. Importantly, this $A$ is an improving set w.r.t. the original labeling, before doing any relabeling of the sequence. If, after the sequence, we have an improving set $A$ in some area which we previously already processed in phase $j$, then this lemma implies that we would have already seen an improving set $A^{\prime}$ in this same area in phase $j$. Because $A^{\prime}$ will have a slightly worse improving ratio than $A$, we have to increase the improving ratio $R$ between every phase to make this argument work.
@@
Let $(A_{i},\ell_{i})_{i=1}^{h}$ be a sequence of $\beta$-improving sets of strong diameter at most $c_{3}\operatorname{log}n/\varepsilon$, for some $\beta>0$ and $0<\varepsilon<\beta$, where $c_{2}$ is the constant given by Lemma 4.5. Then, for each $i\in[h]$ and each $v\in A_{i}$, there exists a large enough constant $c_{2}$ and an improving set $(A,\ell_{A})$ of diameter $c_{3}\operatorname{log}n/\varepsilon$, such that $\operatorname{IR}(A,\ell,\ell_{A})\geq\beta-2\varepsilon$, that is fully contained in the set $\mathcal{N}_{c_{3}\operatorname{log}^{2}n/\varepsilon^{2}}\cap(\cup_{i\in[h]}A_{i})$.
The proof of Lemma 4.6 is also quite involved, and we defer it to Section 4.1. We first proceed with the proof of Lemma 4.4 and then with that of Theorem 3.7.
@@
For any phase $i>0$, let $\mathcal{C}^{(i)}_{1},\ldots,\mathcal{C}^{(i)}_{k_{i}}$ be the clusters determined by MPX at phase $i$ (Line 9), and let $\ell_{i}$ be the labeling of $G$ at the end of phase $i$ (Lines-18 and 23). We proceed by induction on the phase $i$.
If $i=1$, $R_{1}=R$ as initialized at Line 4. An improving set $(A,\ell_{A})$ with $\operatorname{IR}(A,\ell_{1},\ell_{A})\geq R_{1}$ must be such that there is no $s\in[i_{k}]$ such that $\mathcal{N}_{2r+1}[A]\subseteq C^{(1)}_{s}$. Hence, $\operatorname{dist}(A,B_{1})\leq 2r$ and we have that $A\subseteq\mathcal{N}_{c_{3}\operatorname{log}n/\varepsilon+2r}[B_{1}]$, as the strong diameter of $A$ is at most $c_{3}\operatorname{log}n/\varepsilon$. Now, assume $i>1$ and the thesis to be true for all phases $j=1,\ldots,i-1$. Let $(A,\ell_{A})$ be an improving set such that $\operatorname{IR}(A,\ell_{i},\ell_{A})\geq R_{i}$. By Lemma 4.5 again, $A\subseteq\mathcal{N}_{c_{3}\operatorname{log}n/\varepsilon+2r}[B_{i}]$, otherwise we are breaking maximality in Line 14. Suppose, for contradiction, that $A$ is not contained in
@@
This implies that there is a $j^{\star}\leq i-1$ such that $A$ is not contained in $\mathcal{N}_{t_{2}}[B_{j}]$. Note that the strong diameter of $A$ is $c_{3}\operatorname{log}n/\varepsilon$ and at least one node of $A$ is not in $\mathcal{N}_{t_{2}}[B_{j}]$. As a result the set $\mathcal{N}_{c_{2}\operatorname{log}^{2}n/\varepsilon^{2}+2r+1}[A]$ is fully contained in some cluster $C^{(j^{\star})}_{s}$, for some $s\in[i_{j^{\star}}]$. Note that, because the $R_{1},R_{2},\ldots$ are strictly increasing, $(A,\ell_{A})$ is also an improving set such that $\operatorname{IR}(A,\ell_{i},\ell_{A})\geq R_{j^{\star}}$.
We will now describe the changes done by our algorithm after phase $j$ as a single improving sequence. Let us consider the following sequence of $R_{j^{\star}+1}$-improving sets. We pick an arbitrary ordering of clustering at phase $j^{\star}\leq m\leq i$ that is $C^{(m)}_{1},\ldots,C^{(m)}_{k_{m}}$. We say that $C^{(m_{1})}_{s_{1}}<C^{(m_{2})}_{s_{2}}$ if and only if:

### v3

@@ 3.2 Algorithm description
3:Initialize $\lambda\leftarrow$ minimum improvement associated to $(\Pi,\Psi)$
@@
5:Initialize $\varepsilon\leftarrow\frac{\lambda}{100c_{1}\operatorname{log}n}$ $\triangleright$ $c_{1}$ is a constant given by the proof of Theorem 3.7
@@
9:  Run MPX with $\rho=10\cdot c_{2}\cdot\varepsilon^{2}/(c\operatorname{log}^{2}n)$ $\triangleright$ $d=O(\operatorname{log}n/\rho)$ by Lemma 3.6
@@
We now give a formal description of the algorithm given in Algorithm 1. It solves an LOP $(\Pi,\Psi)$ as defined in Definition 2.7. It takes as input the number of nodes $n$ and the description of the LOP $(\Pi,\Psi)$. In more detail, it gets the full description of $\Pi=(\mathcal{V}_{\operatorname{in}},\mathcal{V}_{\operatorname{out}},\mathcal{E}_{\operatorname{in}},\mathcal{E}_{\operatorname{out}},\mathcal{C})$, where $\mathcal{C}$ is an $(2r,\Delta)$-set of constraints over $(\mathcal{V}_{\operatorname{in}}\times\mathcal{V}_{\operatorname{out}},\mathcal{E}_{\operatorname{in}}\times\mathcal{E}_{\operatorname{out}})$, and the potential function $\Psi$ that assigns a non-negative real value to every element of $\mathcal{L}(\mathcal{V}_{\operatorname{in}}\times\mathcal{V}_{\operatorname{out}},\mathcal{E}_{\operatorname{in}}\times\mathcal{E}_{\operatorname{out}},r,\Delta)$.
@@
For every labeled centered graph $(H,v_{H})\in\mathcal{L}(\mathcal{V}_{\operatorname{in}}\times\mathcal{V}_{\operatorname{out}},\mathcal{E}_{\operatorname{in}}\times\mathcal{E}_{\operatorname{out}},2r,\Delta)$ that does not belong to $\mathcal{C}$, there is a way to reassign the labels to $v_{H}$ and its incident half-edges so that the potential $\operatorname{Pot}(G,\ell)$ decreases and the centered graph $(H,v_{H})$ with the new labeling belongs to $\mathcal{C}$. Let $\ell^{\prime}$ be the new labeling: $(\{v_{H}\},\ell^{\prime})$ is a minimal improving set with some positive improvement $\hat{\lambda}$. Let $\lambda$ be the minimum over all possible values of $\hat{\lambda}$. We say that $\lambda$ is the minimum improvement associated to $(\Pi,\Psi)$ (which is initialized in Line 3). Similarly, let $\Lambda$ be the maximum possible decrease of the potential over any centered graph $(H,v_{H})\in\mathcal{L}(\mathcal{V}_{\operatorname{in}}\times\mathcal{V}_{\operatorname{out}},\mathcal{E}_{\operatorname{in}}\times\mathcal{E}_{\operatorname{out}},2r,\Delta)$. We say that $\Lambda$ is the maximum improvement associated to $(\Pi,\Psi)$.
@@
The idea of the algorithm is as follows. Set $\varepsilon=\lambda/(100c_{1}\operatorname{log}(n))$, where $c_{1}$ is a small enough constant (Line 5), and set a threshold improving ratio to be $R=\lambda/4$ (Line 4). We start with an arbitrary labeling of the nodes and half-edges of $G$. Then, we have $\Theta(\operatorname{log}n)$ phases (from Line 8 to Line 26), where in each phase we run the MPX clustering algorithm (Line 9) with parameters $\rho=10\cdot c_{2}\cdot\varepsilon^{2}/(c\operatorname{log}^{2}n)=O(\operatorname{log}^{4}n)$ and $d=O(\operatorname{log}n/\rho)$, where $c,c_{2}$ are large enough constants that come from Lemmas 3.6 and 4.8, respectively: the bound on $d$ then holds by Lemma 3.6. Each run of MPX takes $O(d)=O(\operatorname{log}^{5}n)$ rounds. We assume that each cluster $C$ given by the MPX algorithm comes with a leader node $v_{C}$, that can be the center of the cluster itself. This anyway can be obtained in $O(d)$ rounds after each MPX run. Each leader node looks at the subgraph $H$ of $G$ induced by the cluster $C$ it belongs to, and it computes a maximal sequence of $R$-improving sets $(A_{i},\ell_{i})_{i=1}^{k}$ in $H$ w.r.t. $\ell$. The sequence satisfies that $\mathcal{N}_{2r+1}[A_{i}]\subseteq C$ for all $i\in[k]$ (from Line 13 to Line 14), ensuring that the sequences of different clusters do not interfere with each other. Furthermore, we require the diameter of each set in the sequence to be at most $c_{3}\operatorname{log}n/\varepsilon$ (the choice of the large constant $c_{3}$ comes from Lemma 4.5). When we say that the sequence is maximal, we mean that in $H$ labeled by $\ell_{k}$ there is no minimal improving set $(A^{\prime},\ell^{\prime})$ such that $\operatorname{IR}(A^{\prime},\ell_{k},\ell^{\prime})>R$ and $\mathcal{N}_{2r+1}[A^{\prime}]\subseteq H$. Note that such a sequence must exist and terminate, as the potential of $H$ is always reduced after every relabeling (and cannot be reduced arbitrarily): furthermore, the sequence can be computed in $O(d)$ rounds by brute-force. For this the leader simply collects the topology and labeling of its entire cluster and brute forces every possible sequence in local computation. Then, $v_{C}$ adopts the labeling defined by $\ell_{k}$ and also broadcasts $\ell_{k}$ to every node of the cluster $C$ (Line 17). Every non-leader node just waits for the leader node to broadcast the new labeling (lines 20 and 21), and then it adopts the new labeling (Line 22). Then, the phase terminates and the nodes update the reference value of the improving ratio $R$ to $R+\lambda/(20c_{1}\operatorname{log}n)$ (Line 25). The algorithm terminates after $\Theta(\operatorname{log}n)$ phases, and it returns the labels of the nodes and of the incident half-edges (Line 27).
@@ 4 Analysis of the algorithm
For the rest of the section, we fix an LOP of interest $(\Pi,\Psi)$ and any input graph $G=(V,E)$ of $n$ nodes that is already $(\mathcal{V}_{\operatorname{in}}\times\mathcal{V}_{\operatorname{out}},\mathcal{E}_{\operatorname{in}}\times\mathcal{E}_{\operatorname{out}})$-labeled (hence it has an input and a – possibly invalid – output labeling). Let $\ell$ denote the labeling of $G$. Recall that $\lambda$ is the minimum improvement associated to $(\Pi,\Psi)$, while $\Lambda$ is the maximum improvement associated to $(\Pi,\Psi)$. Trivially, $\lambda\leq\Lambda$. We remark that all the values $r,\Delta,\lambda,\Lambda$ are constant values and will be hidden in the asymptotic notation. In the proofs of the lemmas in Section 4.1, we do keep track of a loose dependency in these parameters, which is polynomial. The reader might extract a $\text{poly}(\Delta^{r},\Lambda)\cdot\text{polylog}(n)$-time algorithm from this analysis; however, we remark that we did not optimize such dependencies.
We fix some standard notation.
@@
For any subset of nodes $A\subseteq V$, we denote by $\ell^{\star}_{A}$ the best possible relabeling of the nodes of $A$ and their incident half-edges, that is,
$\ell^{\star}_{A}=\arg\max_{x\in\mathscr{L}}\set{\operatorname{Imp}(A,\ell,x)}$,
breaking ties arbitrarily, where $\mathscr{L}$ is the set of all possible relabelings of $G$ such that $G$ is a $(\mathcal{V}_{\operatorname{in}}\times\mathcal{V}_{\operatorname{out}},\mathcal{E}_{\operatorname{in}}\times\mathcal{E}_{\operatorname{out}})$-labeled graph where the input labels are untouched, and the output labels might change only on the nodes of $A$ and their incident half-edges.
The whole analysis of the algorithm relies on the following two lemmas (Lemmas 4.2 and 4.4), the last of which requires involved technical claims. The first lemma states that each phases terminates.
@@
For any phase $j$ of the algorithm, let $\ell_{j}$ be the labeling of the algorithm after finishing phase j1j-1 (Lines 1823) and $R_{j}$ be the value of $R$ at the end of phase j1j-1 (Line 25). Furthermore, let $C^{(j)}_{1},\ldots,C^{(j)}_{k}$ be the clusters that are the result of the MPX algorithm at phase $j$ (Line 9), Then, for each cluster $C^{(j)}_{i}$, any maximal sequence of $R_{j}$-improving sets is finite.
@@
So $B_{i}$ is the set of all nodes that have an edge that is crossing different clusters after running MPX at phase $i$. The next lemma shows that after phase $i$, all remaining improving sets of small diameter that have improving ratio at least $R_{i}$ (the value of $R$ initialized at the end of phase i1i-1 at Line 25) are close to all borders $B_{j}$ for all $j\leq i$. Note that after the last phase the intersection of all border nodes will be empty with high probability. So this lemma implies that at the end of our algorithm, there do not exist any improving sets with small diameter and good improving ratio anymore.
@@
At the end of phase $i$ of the algorithm, all minimal improving sets of diameter at most $c_{3}\operatorname{log}n/\varepsilon$ and improving ratio at least $R_{i}$ (the value of $R$ initialized at the end of phase i1i-1 at Line 25) are fully contained in the set
@@
The proof of Lemma 4.4 relies on the following two lemmas. The first one states that for any minimal improving set $A$ of some improving ratio $\beta$, for any node $v\in A$, we can find, within some radius-$O(\operatorname{log}n/\varepsilon)$ neighborhood of $v$, a minimal improving set $A^{\prime}\subseteq A$ with improving ratio at least $\beta-\varepsilon$. The proof of Lemma 4.5 is involved, and we defer it to Section 4.1.
@@
Let $\ell$ be a labeling of $G=(V,E)$ and let $(S,\ell_{S})$ be a $\beta$-minimal improving set, for some $\beta>0$. Recall that $\lambda$ is the minimum improvement associated to $(\Pi,\Psi)$. Then, for every $0<\varepsilon<\min{\{\beta,\lambda\}}$ and any $v\in S$ there exists a large enough constant $c_{3}$, a $k\leq c_{3}\operatorname{log}n/\varepsilon$, and a subset $A\subseteq S\cap\mathcal{N}_{k}[v]$, such that $\operatorname{IR}(A,\ell,\ell_{A}^{\star})\geq\beta-\varepsilon$. Furthermore, $(A,\ell_{A}^{\star})$ is minimal.
The second lemma states that in a sequence of $\beta$-improving sets with small diameter, before any relabeling, around each node of the sequence there is a minimal improving set $A$ with improving ratio slightly less than $\beta$. Importantly, this $A$ is an improving set w.r.t. the original labeling, before doing any relabeling of the sequence. If, after the sequence, we have an improving set $A$ in some area which we previously already processed in phase $j$, then this lemma implies that we would have already seen a minimal improving set $A^{\prime}$ in this same area in phase $j$. Because $A^{\prime}$ will have a slightly worse improving ratio than $A$, we have to increase the improving ratio $R$ between every phase to make this argument useful. The proof of Lemma 4.6 is also quite involved, and we defer it to Section 4.1.
@@
Let $(A_{i},\ell_{i})_{i=1}^{h}$ be a sequence of $\beta$-improving sets of diameter at most $c_{3}\operatorname{log}n/\varepsilon$, for some $\beta>0$ and $0<\varepsilon<\min{\{\beta,\lambda\}}$, where $c_{3}$ is the constant given by Lemma 4.5, and $\lambda$ is the minimum improvement associated to $(\Pi,\Psi)$. Then, for each $i\in[h]$ and each $v\in A_{i}$, there exists a large enough constant $c_{2}$ and an improving set $(A,\ell_{A})$ of diameter $c_{3}\operatorname{log}n/\varepsilon$, such that $\operatorname{IR}(A,\ell,\ell_{A})\geq\beta-2\varepsilon$, that is fully contained in the set $\mathcal{N}_{c_{2}\operatorname{log}^{2}n/\varepsilon^{2}}[v]\cap(\cup_{j\in[h]}A_{j})$. Furthermore, $(A,\ell_{A})$ is minimal.
We first proceed with the proof of Lemma 4.4 and then with that of Theorem 3.7.
@@
For any phase $i>0$, let $\mathcal{C}^{(i)}_{1},\ldots,\mathcal{C}^{(i)}_{h_{i}}$ be the clusters determined by MPX at phase $i$ (Line 9), let $\ell_{i}$ be the labeling of $G$ at the end of phase $i$ (Lines-18 and 23), and let $R_{i}$ be the improving ratio initialized at the end of phase i1i-1 (Line 25), with $R_{1}=R$ as in Line 4. We proceed by induction on the phase $i$.
If $i=1$, $R_{1}=R$ as initialized at Line 4. A minimal improving set $(A,\ell_{A})$ of diameter at most $c_{3}\operatorname{log}n/\varepsilon$ with $\operatorname{IR}(A,\ell_{1},\ell_{A})\geq R_{1}$ must be such that there is no $s\in[h_{1}]$ such that $\mathcal{N}_{2r+1}[A]\subseteq C^{(1)}_{s}$, otherwise we are breaking maximality in Line 14. Hence, $\operatorname{dist}(A,B_{1})\leq 2r$ and we have that $A\subseteq\mathcal{N}_{c_{3}\operatorname{log}n/\varepsilon+2r}[B_{1}]$, as the diameter of $A$ is at most $c_{3}\operatorname{log}n/\varepsilon$. Now, assume $i>1$ and the statement to be true for all phases $j=1,\ldots,i-1$. Let $(A,\ell_{A})$ be a minimal improving set of diameter at most $c_{3}\operatorname{log}n/\varepsilon$ such that $\operatorname{IR}(A,\ell_{i},\ell_{A})\geq R_{i}$. For the same reason as in the case $i=1$, it holds that $A\subseteq\mathcal{N}_{c_{3}\operatorname{log}n/\varepsilon+2r}[B_{i}]$. Suppose, by contradiction, that $A$ is not contained in
@@
This implies that there is a $j^{\star}\leq i-1$ such that $A$ is not contained in $\mathcal{N}_{t_{2}}[B_{j^{\star}}]$. Note that the diameter of $A$ is at most $c_{3}\operatorname{log}n/\varepsilon$ and at least one node of $A$ is not in $\mathcal{N}_{t_{2}}[B_{j^{\star}}]$. As a result, the set $\mathcal{N}_{c_{2}\operatorname{log}^{2}n/\varepsilon^{2}+2r+1}[A]$ is fully contained in some cluster $C^{(j^{\star})}_{s}$, for some $s\in[h_{j^{\star}}]$.
We can now describe the family of all improving sets that are actually used by our algorithm after phase $j^{\star}$ (Line 14) as a single improving sequence. We will define an ordering of all minimal improving sets that our algorithm has flipped. In order to do that, we first define an ordering of all clusters. We say that $C^{(m_{1})}_{s_{1}}<C^{(m_{2})}_{s_{2}}$ if and only if: